Estimate the number of base pairs in a relaxed plasmid with 86.6 nm diameter. Assume B-form
Answer #1

Ans:- a relaxed plasmid

Diameter = 86.6nm

Radius = D/2 = 43.3nm

We want to estimate the number of base pair in it..

As we know that:-

1base pair = 0.34 nm or 10 bp = 3.4nm

And.

Circumference of plasmid (~ length of linear DNA.)= 2πr

Where π=3.14.

Therefore;

Circumference of plasmid= 2×3.14×43.3nm

= 272 nm(approx)

And

Total base pair in plasmid = circumference of plasmid / distance between one base pair.

Total bp in plasmid = 272/0.34

= 800bp.